50kN/m B 6m — 6m -

i) state clearly the shear force and bending moment equations and

ii) draw the shear force and bending moment diagrams.
Answer #1

50kN/m 6 MB. - Rox 4 ? o 6m 6m RBY here Røy = (x6 x 50 + 50 x 6 )KN = (180 +300) KN = 450KN & RB x=0 MB - fiso x8 + $0 x6 x 3-6m 6m- + Straight line 450 KN Curve 150KN SFD BMD 2100 200 is a reaction Roy from fixed end. Shearforce Equation let us takeLxXx 50 = 25%. 1 12<x<61 from A side Bending Moment Equation, M --2100 + 50.X• ķ = - 2100 + 25x2 [0<x<6] from Bride M = £ x 5